OpenStudy (anonymous):

Find the laplace transform of:

OpenStudy (anonymous):

\[10\sin \frac{ t }{ 2 }\]

OpenStudy (anonymous):

isnt it just \[10\left( \frac{ 1/2 }{ s ^{2}+\left( \frac{ 1 }{ 2 } \right)^{2} } \right)\]

OpenStudy (anonymous):

Or do you want it by definition?

OpenStudy (anonymous):

it is a t not 1

OpenStudy (anonymous):

I know, the t goes away after the transform, isnt that the point?

OpenStudy (anonymous):

well ya, but it just vanishes?

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/pdf/Laplace_Table.pdf

OpenStudy (anonymous):

see #7

OpenStudy (anonymous):

ah, Thanks